# Self-similar sets and Lipschitz graphs

### v1

@@ 7.3. Ergodic theory
There exists $\theta\in[0,\pi)$ so that for every $n\in\mathbb{N}$, $D(n;\theta)\in\left[1-\frac{\varepsilon}{2},1\right]$
@@ 7.4. Extracting a substantial subset and building the graph
Let $\left\{f_{k}\right\}_{k=1}^{M}$, where each $f_{k}$ is of the form (11), be a $\nu$-non-degenerate IFS that satisfies the open set condition with similarity dimension $1$. Let $H$ be the attractor of $\left\{f_{k}\right\}_{k=1}^{M}$, let $a$ and $b$ denote the Ahlfors lower and upper constants of $H$, respectively, let $\omega$ denote the overlapping index of $\left\{f_{k}\right\}_{k=1}^{M}$, and set $K=\operatorname{conv}(H)$.
There exists a constant $\varepsilon_{0}\left(r_{1},\ldots,r_{M},M,a,b,\omega,K\right)\in(0,1)$ such that for any $\varepsilon\in(0,\varepsilon_{0}]$, there exists $\kappa\in\mathbb{N}$ and a UIFS $\left\{\varphi_{j}\right\}_{j=1}^{N}\subseteq\left\{f_{k^{(\kappa)}}\right\}$ with scale factor
@@
    $E_{n}=\bigsqcup_{i=1}^{M_{n}}K_{i}^{n}$, where each $K_{i}^{n}=\varphi_{j^{(n)}}(K)$ for some $j^{(n)}\in\{1,...,N\}^{n}$.
@@
$\displaystyle c_{4}:=\left(\frac{a\nu}{8\omega^{2}b\alpha_{0}^{2}\operatorname{diam}(K)}\right)^{2}\frac{1}{3c_{e}\left(\alpha_{0}+1\right)}\inf_{\gamma\in\left[0,1\right]}\left[\frac{\operatorname{diam}(K)}{4\alpha_{0}+2}\right]^{\gamma}\frac{1-e^{-(1-\gamma)}}{1-\gamma}>0,$
@@
We now construct the nested sequence of sets. Lemma 3.10 implies that $\left\{\varphi_{j}\right\}_{j=1}^{N}$ inherits the open set condition from $\left\{f_{k}\right\}_{k=1}^{M}$. An application of Lemma 7.3 shows that $\left\{\varphi_{j}\right\}_{j=1}^{N}$ has upper Ahlfors regularity constant
(45) $b_{0}=\frac{8\omega b\operatorname{diam}(K)}{a\nu}r^{\gamma-1}$
@@
Lemma 3.14 implies that the generations $\left\{C_{n}\right\}_{n=0}^{\infty}$ have overlapping index $\omega_{0}\leq\omega$. Therefore, we may apply Lemma 7.7 to $\left\{\varphi_{j}\right\}_{j=1}^{N}$ and $K$ with $n=1$ to deduce that there exists a constant $\delta_{0}>0$ and a set $J\subseteq[0,\pi)$ such that $\left|[0,\pi)\backslash J\right|\leq\varepsilon$ and the following holds:
(46) For every $\phi\in J$, there exists a subset $C_{1,\phi}\subseteq C_{1}$ with $N_{\phi}\geq\varepsilon\delta_{0}r^{-\gamma}$
connected components for which $P_{\phi}(C_{1,\phi})$ is a disjoint union of $N_{\phi}$
$r$-separated intervals.
@@
$N_{\phi}$ $\geq\varepsilon\delta_{0}r^{-\gamma}=\frac{\mathcal{H}^{\gamma}(C)}{15c_{e}\alpha_{0}^{4}\left(\alpha_{0}+1\right)\omega_{0}^{2}b_{0}}\frac{1-e^{-(1-\gamma)}}{(2+4\alpha_{0})^{\gamma}}\varepsilon r^{-\gamma}$
$\geq\left(\frac{a\nu}{8\omega^{2}b\alpha_{0}^{2}\operatorname{diam}(K)}\right)^{2}\frac{1}{3c_{e}\left(\alpha_{0}+1\right)}\left[\frac{\operatorname{diam}(K)}{2+4\alpha_{0}}\right]^{\gamma}\left(1-e^{-(1-\gamma)}\right)2\eta r^{2-3\gamma}$
@@
where we have applied (45), (37), and $\omega_{0}\leq\omega$, the definition of $c_{4}$, the bounds on $\gamma$, the definition of $\eta_{1}$, and that $\eta\leq\eta_{1}$. From (31), we see that $\frac{3c_{2}}{2}\eta\geq r^{\frac{2\eta}{M}}$ and since $M\geq 3$, we get
@@
Let $A$ denote the rotation matrix for $\left\{\varphi_{j}\right\}_{j=1}^{N}$. The matrix $A$ induces a rotation map on $[0,\pi)$, $T_{A}$, in the sense of Lemma 7.9. For each $\theta\in[0,\pi)$, define $\mathcal{S}(\theta):=\left\{n\in\mathbb{N}~{}:~{}T_{A}^{n-1}(\theta)\in J\right\}$. Since $\left|J\right|\geq\pi-\varepsilon$, then Lemma 7.9 provides an angle $\theta\in[0,\pi)$ such that for every $n\in\mathbb{N}$, it holds that
(48) #$\#(\mathcal{S(\theta)}\cap\left\{1,\ldots,n\right\})\geq\left(1-\frac{\varepsilon}{2}\right)n.$
@@
If $n\in\mathcal{S}(\theta)$, then $P_{\theta}\circ A^{n-1}=P_{T^{n-1}_{A}(\theta)}$ and $\phi_{n}:=T_{A}^{n-1}(\theta)\in J$. Therefore, by (46), there exists $C_{1}(n):=C_{1,\phi_{n}}\subseteq C_{1}$ with $N_{n}:=N_{\phi_{n}}$ connected components for which $P_{\phi_{n}}(C_{1,\phi_{n}})=P_{\theta}(A^{n-1}(C_{1}(n)))$ is a disjoint union of $N_{n}$ $r$-separated intervals. The bound (47) shows that $N_{n}\geq r^{-1+\frac{\varepsilon}{2}}$. We may write
@@
for some $\mathcal{I}_{n}\subseteq\{1,...,N\}$ with #$\#{\mathcal{I}_{n}}=N_{n}$.
If $n\in\mathbb{N}\setminus\mathcal{S}(\theta)$, then we choose $C_{1}(n)\subseteq C_{1}$ to have $N_{n}$ connected components for which the projection $P_{\theta}(A^{n-1}(C_{1}(n)))$ is a disjoint union of $N_{n}$ $r$-separated intervals. Since $n\notin\mathcal{S}(\theta)$, then $T_{A}^{n-1}(\theta)\notin J$, so we cannot guarantee that $N_{n}$ has a large lower bound, but we can ensure that $N_{n}\geq 1$. Again in this case, there is a non-empty subset of indices $\mathcal{I}_{n}\subseteq\{1,...,N\}$, where #$\#{\mathcal{I}_{n}}=N_{n}$, such that (49) holds.
@@
Let $\left\{f_{k}\right\}_{k=1}^{M}$, where each $f_{k}$ is of the form (11), be a $\nu$-non-degenerate IFS that satisfies the open set condition with similarity dimension $1$. Let $H$ be the attractor of $\left\{f_{k}\right\}_{k=1}^{M}$, let $a$ and $b$ denote the Ahlfors lower and upper constants of $H$, respectively, let $\omega$ denote the overlapping index of $\left\{f_{k}\right\}_{k=1}^{M}$, and set $K=\operatorname{conv}(H)$.
There exists $\varepsilon_{0}=\varepsilon_{0}\left(r_{1},\ldots,r_{M},M,a,b,\omega,K\right)\in(0,1)$ so that for any $\varepsilon\in\left(0,\varepsilon_{0}\right]$, there exists a Lipschitz graph $\Gamma$ for which
@@
We check that $\left\{E_{n}\right\}_{n=1}^{\infty}$ satisfies the hypotheses of Proposition 4.1. By construction, $\left\{E_{n}\right\}_{n=1}^{\infty}$ is nested and from item 1, each $E_{n}:=\bigsqcup_{i=1}^{M_{n}}K^{n}_{i}$, where every connected component $K_{i}^{n}=\varphi_{j^{(n)}}(K)$ is a translated and rescaled copy of $K$, hence closed and convex. It follows that $r^{n}\nu\leq\left|P_{\theta}\left(K_{i}^{n}\right)\right|\leq r^{n}\operatorname{diam}(K)$, and then (18) holds with $\lambda=\frac{\operatorname{diam}(K)}{\nu}$. For each $n,i$, $\operatorname{diam}\left(K_{i}^{n}\right)\leq\operatorname{diam}(K){r}^{n}$ and then (20) holds with $c=\operatorname{diam}(K)$, $\sigma=r<1$.
@@ Appendix A The Measure of the 4-corner Cantor Set
Let $\mathcal{C}_{4}$ denote the 4-corner Cantor set. Since projections decrease measure, then we know that $\mathcal{H}^{1}(\mathcal{C}_{4})\geq\mathcal{H}^{1}\left(P_{\arctan(1/2)}\left(\mathcal{C}_{4}\right)\right)=\frac{3}{\sqrt{5}}$. See Figure 7. On the other hand, if we choose the cubes involved in the construction of $\mathcal{C}_{4}$ as the covering in the definition of Hausdorff measure, then it follows that $\mathcal{H}^{1}(\mathcal{C}_{4})\leq\sqrt{2}$. Therefore, $\mathcal{H}^{1}(\mathcal{C}_{4})\in\left[\frac{3}{\sqrt{5}},\sqrt{2}\right]$.
We include a computation of the exact Hausdorff measure of $\mathcal{C}_{4}$. This result, originally due to Davies, was communicated in private communication to the second author by Kenneth Falconer. To our knowledge, the only other, less elementary proof of this fact follows from [Mar79].

### v2

@@ 7.3. Ergodic theory
There exists $\theta\in[0,\pi)$ so that for every $n\in\mathbb{N}$, $D(n;\theta)\in\left[1-\frac{\varepsilon}{2},1\right]$ .
@@ 7.4. Extracting a substantial subset and building the graph
Let $\left\{f_{k}\right\}_{k=1}^{M}$, where each $f_{k}$ is of the form (11), be a $\nu$-non-degenerate IFS that satisfies the open set condition with similarity dimension $1$. Let $H$ be the attractor of $\left\{f_{k}\right\}_{k=1}^{M}$, let $a$ and $b$ denote the Ahlfors lower and upper constants of $H$, respectively, let $\tau$ denote the overlapping index of $\left\{f_{k}\right\}_{k=1}^{M}$, and set $K=\operatorname{conv}(H)$.
There exists a constant $\varepsilon_{0}\left(r_{1},\ldots,r_{M},M,a,b,\tau,K\right)\in(0,1)$ such that for any $\varepsilon\in(0,\varepsilon_{0}]$, there exists $\kappa\in\mathbb{N}$ and a UIFS $\left\{\varphi_{j}\right\}_{j=1}^{N}\subseteq\left\{f_{k^{(\kappa)}}\right\}$ with scale factor
@@
    $E_{n}=\bigcup_{i=1}^{M_{n}}K_{i}^{n}$, where each $K_{i}^{n}=\varphi_{j^{(n)}}(K)$ for some $j^{(n)}\in\{1,...,N\}^{n}$.
@@
$\displaystyle c_{4}:=\left(\frac{a\nu}{8\tau^{2}b\alpha_{0}^{2}\operatorname{diam}(K)}\right)^{2}\frac{1}{3c_{e}\left(\alpha_{0}+1\right)}\inf_{\gamma\in\left[0,1\right]}\left[\frac{\operatorname{diam}(K)}{4\alpha_{0}+2}\right]^{\gamma}\frac{1-e^{-(1-\gamma)}}{1-\gamma}>0,$
@@
We now construct the nested sequence of sets. Lemma 3.10 implies that $\left\{\varphi_{j}\right\}_{j=1}^{N}$ inherits the open set condition from $\left\{f_{k}\right\}_{k=1}^{M}$. An application of Lemma 7.3 with $\dim(H)=1$ shows that $\left\{\varphi_{j}\right\}_{j=1}^{N}$ has upper Ahlfors regularity constant
(46) $b_{0}=\frac{8\tau b\operatorname{diam}(K)}{a\nu}r^{\gamma-1}$
@@
Lemma 3.14 implies that the generations $\left\{C_{n}\right\}_{n=0}^{\infty}$ have overlapping index $\tau_{0}\leq\tau$. Therefore, we may apply Lemma 7.7 to $\left\{\varphi_{j}\right\}_{j=1}^{N}$ and $K$ with $n=1$ to deduce that there exists a constant $\delta_{0}>0$ and a set $J\subseteq[0,\pi)$ such that $\left|[0,\pi)\backslash J\right|\leq\varepsilon$ and the following holds:
(47) For every $\phi\in J$, there exists a subset $C_{1,\phi}\subseteq C_{1}$ with $N_{\phi}\geq\varepsilon\delta_{0}r^{-\gamma}$ connected
components for which $P_{\phi}(C_{1,\phi})$ is a union of $N_{\phi}$ $r$-separated intervals.
@@
$N_{\phi}$ $\geq\varepsilon\delta_{0}r^{-\gamma}=\frac{\mathcal{H}^{\gamma}(C)}{15c_{e}\alpha_{0}^{4}\left(\alpha_{0}+1\right)\tau_{0}^{2}b_{0}}\frac{1-e^{-(1-\gamma)}}{(2+4\alpha_{0})^{\gamma}}\varepsilon r^{-\gamma}$
$\geq\left(\frac{a\nu}{8\tau^{2}b\alpha_{0}^{2}\operatorname{diam}(K)}\right)^{2}\frac{1}{3c_{e}\left(\alpha_{0}+1\right)}\left[\frac{\operatorname{diam}(K)}{2+4\alpha_{0}}\right]^{\gamma}\left(1-e^{-(1-\gamma)}\right)2\eta r^{2-3\gamma}$
@@
where we have applied (46), (38), and $\tau_{0}\leq\tau$, the definition of $c_{4}$, the bounds on $\gamma$, the definition of $\eta_{1}$, and that $\eta\leq\eta_{1}$. From (32), we see that $\frac{3c_{2}}{2}\eta\geq r^{\frac{2\eta}{M}}$ and since $M\geq 3$, we get
@@
Let $A$ denote the rotation matrix for $\left\{\varphi_{j}\right\}_{j=1}^{N}$. The matrix $A$ induces a rotation map on $0,\pi)$, $T_{A}$, in the sense of Lemma 7.9. For each $\theta\in[0,\pi)$, define $\mathcal{S}(\theta):=\left\{n\in\mathbb{N}~:~T_{A}^{n-1}(\theta)\in J\right\}$. Since $\left|J\right|\geq\pi-\varepsilon$, then Lemma 7.9 provides an angle $\theta\in[0,\pi)$ such that for every $n\in\mathbb{N}$, it holds that
(49) $(\mathcal{S(\theta)}\cap\left\{1,\ldots,n\right\})\geq\left(1-\frac{\varepsilon}{2}\right)n.$
@@
If $n\in\mathcal{S}(\theta)$, then $P_{\theta}\circ A^{n-1}=P_{T^{n-1}_{A}(\theta)}$ and $\phi_{n}:=T_{A}^{n-1}(\theta)\in J$. Therefore, by (47), there exists $C_{1}(n):=C_{1,\phi_{n}}\subseteq C_{1}$ with $N_{n}:=N_{\phi_{n}}$ connected components for which $P_{\phi_{n}}(C_{1,\phi_{n}})=P_{\theta}(A^{n-1}(C_{1}(n)))$ is a union of $N_{n}$ $r$-separated intervals. The bound (48) shows that $N_{n}\geq r^{-1+\frac{\varepsilon}{2}}$. We may write
@@
for some $\mathcal{I}_{n}\subseteq\{1,...,N\}$ with ${\mathcal{I}_{n}}=N_{n}$.
If $n\in\mathbb{N}\setminus\mathcal{S}(\theta)$, then we choose $C_{1}(n)\subseteq C_{1}$ to have $N_{n}$ connected components for which the projection $P_{\theta}(A^{n-1}(C_{1}(n)))$ is a union of $N_{n}$ $r$-separated intervals. Since $n\notin\mathcal{S}(\theta)$, then $T_{A}^{n-1}(\theta)\notin J$, so we cannot guarantee that $N_{n}$ has a large lower bound, but we can ensure that $N_{n}\geq 1$. Again in this case, there is a non-empty subset of indices $\mathcal{I}_{n}\subseteq\{1,...,N\}$, where ${\mathcal{I}_{n}}=N_{n}$, such that (50) holds.
@@
Let $\left\{f_{k}\right\}_{k=1}^{M}$, where each $f_{k}$ is of the form (11), be a $\nu$-non-degenerate IFS that satisfies the open set condition with similarity dimension $1$. Let $H$ be the attractor of $\left\{f_{k}\right\}_{k=1}^{M}$, let $a$ and $b$ denote the Ahlfors lower and upper constants of $H$, respectively, let $\tau$ denote the overlapping index of $\left\{f_{k}\right\}_{k=1}^{M}$, and set $K=\operatorname{conv}(H)$.
There exists $\varepsilon_{0}=\varepsilon_{0}\left(r_{1},\ldots,r_{M},M,a,b,\tau,K\right)\in(0,1)$ so that for any $\varepsilon\in\left(0,\varepsilon_{0}\right]$, there exists a Lipschitz graph $\Gamma$ for which
@@
We check that $\left\{E_{n}\right\}_{n=1}^{\infty}$ satisfies the hypotheses of Proposition 4.1. By construction, $\left\{E_{n}\right\}_{n=1}^{\infty}$ is nested and from item 1, each $E_{n}:=\bigcup_{i=1}^{M_{n}}K^{n}_{i}$, where every connected component $K_{i}^{n}=\varphi_{j^{(n)}}(K)$ is a translated and rescaled copy of $K$, hence closed and convex. It follows that $r^{n}\nu\leq\left|P_{\theta}\left(K_{i}^{n}\right)\right|\leq r^{n}\operatorname{diam}(K)$, and then (18) holds with $\lambda=\frac{\operatorname{diam}(K)}{\nu}$. For each $n,i$, $\operatorname{diam}\left(K_{i}^{n}\right)\leq\operatorname{diam}(K){r}^{n}$ and then (20) holds with $\sigma_{n}=\operatorname{diam}(K){r}^{n}$.
@@ Appendix A The Measure of the 4-corner Cantor Set
Let $\mathcal{C}_{4}$ denote the 4-corner Cantor set. Since projections decrease measure, we know that $\mathcal{H}^{1}(\mathcal{C}_{4})\geq\mathcal{H}^{1}\left(P_{\arctan(1/2)}\left(\mathcal{C}_{4}\right)\right)=\frac{3}{\sqrt{5}}$. See Figure 7. On the other hand, if we choose the cubes involved in the construction of $\mathcal{C}_{4}$ as the covering in the definition of Hausdorff measure, then it follows that $\mathcal{H}^{1}(\mathcal{C}_{4})\leq\sqrt{2}$. Therefore, $\mathcal{H}^{1}(\mathcal{C}_{4})\in\left[\frac{3}{\sqrt{5}},\sqrt{2}\right]$.
We include a computation of the exact Hausdorff measure of $\mathcal{C}_{4}$. This result, originally due to Davies, was communicated in private communication to the second author by Kenneth Falconer. Other, less elementary proofs of this fact can be found in [Mar79] and [ZW99].